OpenStudy (anonymous):

Stephanie finances $423,000 with a 20/5 balloon mortgage at 4.45%. How much will she pay for principal and interest over the life of the loan? A. $508,077.56 B. $509,378.67 C. $ 510,674.71 D. $505,422.01

OpenStudy (anonymous):

Determine the Payments for a fixed morgage of 20 years, balance it out for 5 years, and pay off the remaining balance. add up all the payments made

OpenStudy (anonymous):

a fixed loan can be modeled using this formula: \[A=-Bk^n+P\frac{1-k^n}{1-k}\] When A = 0 the loan is paid in full \[0=-Bk^n+P\frac{1-k^n}{1-k}\] \[Bk^n=P\frac{1-k^n}{1-k}\] \[Bk^n\frac{1-k}{1-k^n}=P\] there are 5(12) payments \[60*Bk^n\frac{1-k}{1-k^n}=$paid~in~5~years\] and after 5 years the remianing balance is \[A=-Bk^{60}+P\frac{1-k^{60}}{1-k}\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28423000%29k%5E%2820*12%29%5Cfrac%7B1-k%7D%7B1-k%5E%2820*12%29%7D%3B+k%3D1%2B.0445%2F12 using 20*12 as the number of periods, and k=1+.0445/12 as the compunding interest Payments are about 2664.70 a month, 5 years gives us a total of 159 882, plus the remaining balance of the loan

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-%28423000%29k%5E%7B60%7D%2B%282664.70%29%5Cfrac%7B1-k%5E%7B60%7D%7D%7B1-k%7D%3B+k%3D1%2B.0445%2F12 so there is about 349 497 left to pay off 349 497 159 882 -------- *******